Please help me solve this equation.
6a^2-4=5a

Answer by TakeATuition.com(57)

SOLUTION -
6a^2 - 5a - 12 = 0
6a^2 - 8a + 3a - 12 = 0
2a(3a - 4) + 1(3a - 4) = 0
(3a - 4)(2a + 1) = 0
3a - 4 = 0
3a = 4
a = 4/3
2a + 1 = 0
2a = -1
a = -1/2

FOR MORE HELP VISIT : www.TakeATuition.com